# The algebraic and geometric classification of

### v1

@@ 1.6. The algebraic classification of compatible Novikov algebras
Theorem 12.
Let $\mathcal{C}$ be a nonzero $2$-dimensional compatible Novikov algebra. Then $\mathcal{C}$ is isomorphic to one and only one compatible commutative associative algebra listed in Theorem 8 or one of the following algebras:
$\mathcal{C}_{01}$ :::: $e_{1}\ast e_{1}=e_{1}+e_{2},\ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{04}$ :::: $e_{2}\ast e_{1}=e_{1}.$
$\mathcal{C}_{06}^{\alpha\neq 1}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=\alpha e_{2},\ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{09}^{\alpha,\beta\neq 0}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\\ \end{tabular}\right.$
$\mathcal{C}_{10}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{13}^{\alpha\neq 1}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{14}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{1}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{20}^{\alpha,0}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&\\ $e_{2}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=e_{1}.$\end{tabular}\right.$
$\mathcal{C}_{21}^{\alpha,0}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&\\ $e_{2}\ast e_{1}=\alpha e_{1}.$\end{tabular}\right.$
$\mathcal{C}_{22}^{\alpha\neq 0,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=e_{1}+\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{23}^{\alpha\neq 0,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{31}^{(\alpha,\beta,\gamma)\neq(1,\beta,\beta)}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{32}^{(\alpha,\beta,\gamma)\neq(1,\beta,\beta)}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{33}$ :::: {
=e1e1e1, =e2e1e2,
=e1e2e1, =e2e2e2.
[…]
@@
2. The geometric classification of compatible algebras
2.1. Degenerations and the geometric classification of algebras
@@
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\cong\textrm{Hom}({\mathbb{V}}^{\otimes 2},{\mathbb{V}})\cong({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}}$
@@
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\cong({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}}\oplus({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}}$
@@
$\mu(e_{i},e_{j})=\sum_{p=1}^{n}c_{ij}^{k}e_{k}\textrm{ and }\mu^{\prime}(e_{i},e_{j})=\sum_{p=1}^{n}c_{ij}^{\prime k}e_{k},$
which corresponds to a point in the affine space $\mathbb{C}^{2n^{3}}$. Then a set of bilinear pairs $\mathcal{S}$ corresponds to an algebraic variety, i.e., a Zariski closed set, if there are some polynomial equations in variables $c_{ij}^{k},c_{ij}^{\prime k}$ with zero locus equal to the set of structural constants of the bilinear pairs in $\mathcal{S}$. Given the identities defining a particular class of compatible $\Omega$-algebras, we can obtain a set of polynomial equations in variables $c_{ij}^{k},c_{ij}^{\prime k}$. This class of $n$-dimensional compatible $\Omega$-algebras is a variety. Denote it by $\mathcal{T}_{n}$.Now, consider the following action of $\rm{GL}({\mathbb{V}})$ on ${\mathcal{T}}_{n}$:
$(g*(\mu,\mu^{\prime}))(x,y):=(g\mu(g^{-1}x,g^{-1}y),g\mu^{\prime}(g^{-1}x,g^{-1}y))$
@@
A geometric classification of a variety of algebras consists on describing the irreducible components of the variety. Recall that any affine variety can be represented as a finite union of its irreducible components uniquely. Note that describing the irreducible components of ${\mathcal{T}_{n}}$ gives us the rigid algebras of the variety, which are those bilinear pairs with an open $\textrm{GL}(\mathbb{V})$-orbit. This is due to the fact that a bilinear pair is rigid in a variety if and only if the closure of its orbit is an irreducible component of the variety. For this reason, the following notion is convenient. Denote by $\overline{{\mathcal{O}}((\mu,\mu^{\prime}))}$ the closure of the orbit of $(\mu,\mu^{\prime})\in{\mathcal{T}_{n}}$.
Definition 13.
@@
Furthermore, for a parametric family of algebras we have the following notion.
Definition 14.
@@
Let ${{\rm T}}(*)=\{{{\rm T}}(\alpha):{\alpha\in I}\}$ be a family of $n$-dimensional compatible $\Omega$-algebras of a fixed class corresponding to ${\mathcal{T}}_{n}$ and let ${{\rm T}}^{\prime}$ be another algebra. Suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. We say that the family ${{\rm T}}(*)$ degenerates to ${{\rm T}}^{\prime}$, and write ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{\{{\mathcal{O}}((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$.Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{\{{\mathcal{O}}((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$ then we call it a non-degeneration, and we write ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$.
@@
Observe that ${\rm T}^{\prime}$ corresponds to an irreducible component of $\mathcal{T}_{n}$ (more precisely, $\overline{{\rm T}^{\prime}}$ is an irreducible component) if and only if ${{\rm T}}\not\to{{\rm T}}^{\prime}$ for any $n$-dimensional compatible $\Omega$-algebra ${\rm T}$ and ${{\rm T}(*)}\not\to{{\rm T}}^{\prime}$ for any parametric family of $n$-dimensional compatible $\Omega$-algebras ${\rm T}(*)$. To prove a particular algebra corresponds to an irreducible compotent, we will use the next ideas.Firstly, since $\mathrm{dim}\,{\mathcal{O}}((\mu,\mu^{\prime}))=n^{2}-\mathrm{dim}\,\mathfrak{Der}({\rm T})$, then if ${\rm T}\to{\rm T}^{\prime}$ and ${\rm T}\not\cong{\rm T}^{\prime}$, we have that $\mathrm{dim}\,\mathfrak{Der}({\rm T})<\mathrm{dim}\,\mathfrak{Der}({\rm T}^{\prime})$, where $\mathfrak{Der}({\rm T})$ denotes the Lie algebra of derivations of ${\rm T}$. Secondly, to prove degenerations, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two compatible $\Omega$-algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\mathcal{T}}}_{n}$, respectively. Let $c_{ij}^{k},$ $c_{ij}^{\prime k}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there exist $n^{2}$ maps $a_{i}^{j}(t):\mathbb{C}^{*}\to\mathbb{C}$ such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{ij}^{k}(t),c_{ij}^{\prime k}(t)$ of $(\mu,\mu^{\prime})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{ij}^{k}(t)=c_{ij}^{k}$ and $\lim\limits_{t\to 0}c_{ij}^{\prime k}(t)=c_{ij}^{\prime k}$, then ${{\rm T}}\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ is called a parametrized basis for ${{\rm T}}\to{{\rm T}}^{\prime}$. Thirdly, to prove non-degenerations we may use a remark that follows from this lemma, see [3].
Lemma 15.
@@
Consider two compatible $\Omega$-algebras ${\rm T}$ and ${\rm T}^{\prime}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Let C be a Zariski closed in ${\mathcal{T}}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu,\mu^{\prime})$ of ${\rm T}$ in C, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in C.
@@
Remark 16.
Moreover, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two compatible $\Omega$-algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${\mathcal{T}_{n}}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Then if $\mu,\mu^{\prime},\lambda,\lambda^{\prime}$ represents algebras ${\rm T}_{0},{\rm T}_{1},{\rm T}^{\prime}_{0},{\rm T}^{\prime}_{1}$ in the affine space $\mathbb{C}^{n^{3}}$ of algebras with a single multiplication, respectively, we have ${\rm T}_{0}\to{\rm T}^{\prime}_{0}$ and ${\rm T}_{1}\to{\rm T}^{\prime}_{1}$. So for example, $(0,\mu)$ can not degenerate in $(\lambda,0)$ unless $\lambda=0$.
Fourthly, to prove ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. Let $c_{ij}^{k},c_{ij}^{\prime k}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there is a pair of maps $(f,(a_{i}^{j}))$, where $f:\mathbb{C}^{*}\to I$ and $a_{i}^{j}:\mathbb{C}^{*}\to\mathbb{C}$ are such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{ij}^{k}(t),c_{ij}^{\prime k}(t)$ of $(\mu\big{(}f(t)\big{)},\mu^{\prime}\big{(}f(t)\big{)})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{ij}^{k}(t)=c_{ij}^{k}$ and $\lim\limits_{t\to 0}c_{ij}^{\prime k}(t)=c_{ij}^{\prime k}$, then ${{\rm T}}(*)\to{{\rm T}}^{\prime}$. In this case $E_{1}(t),\dots,E_{n}(t)$ and $f(t)$ are called a parametrized basis and a parametrized index for ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, respectively. Fifthly, to prove ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$, we may use an analogous of Remark 16 for parametric families that follows from Lemma 17.
Lemma 17.
Consider the family of compatible $\Omega$-algebras ${\rm T}(*)$ and the compatible $\Omega$-algebra ${\rm T}^{\prime}$. Suppose ${\rm T}(*)\to{\rm T}^{\prime}$. Let C be a Zariski closed in $\mathcal{T}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu(\alpha),\mu^{\prime}(\alpha))$ of ${\rm T}(\alpha)$ in C for every $\alpha\in I$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in C.
@@
2.2. The geometric classification of compatible commutative associative algebras

### v2

@@ 2.6 The algebraic classification of compatible Novikov algebras
Theorem 2.12.
Let $\mathcal{C}$ be a nonzero $2$-dimensional compatible Novikov algebra. Then $\mathcal{C}$ is isomorphic to one and only one compatible commutative associative algebra listed in Theorem 2.8 or one of the following algebras:
$\mathcal{C}_{01}\colon\ e_{1}\ast e_{1}=e_{1}+e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{04}\colon\ e_{2}\ast e_{1}=e_{1},$
$\mathcal{C}_{06}^{\alpha\neq 1}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{09}^{\alpha,\beta\neq 0}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\beta e_{2},e_{2}\ast e_{1}=\alpha e_{2},$
$\mathcal{C}_{10}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},$
$\mathcal{C}_{13}^{\alpha\neq 1}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},$
$\mathcal{C}_{14}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1}+\alpha e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{20}^{\alpha,0}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{2}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=e_{1},$
$\mathcal{C}_{21}^{\alpha,0}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{2}\ast e_{1}=\alpha e_{1},$
$\mathcal{C}_{22}^{\alpha\neq 0,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{1},\quad e_{2}\ast e_{1}=\beta e_{1},\quad e_{2}\ast e_{2}=e_{1}+\alpha e_{2},$
$\mathcal{C}_{23}^{\alpha\neq 0,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{1},\quad e_{2}\ast e_{1}=\beta e_{1},\quad e_{2}\ast e_{2}=\alpha e_{2},$
$\mathcal{C}_{31}^{(\alpha,\beta,\gamma)\neq(1,\beta,\beta)}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1}+e_{2},\quad e_{1}\ast e_{2}=\gamma e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{32}^{(\alpha,\beta,\gamma)\neq(1,\beta,\beta)}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\gamma e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{33}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
@@
3 The geometric classification of compatible algebras
3.1 Degenerations and the geometric classification of algebras
@@
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\cong\textrm{Hom}\bigl{(}{\mathbb{V}}^{\otimes 2},{\mathbb{V}}\bigr{)}\cong({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}}$
@@
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\cong({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}}\oplus({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}},$
@@
$\mu(e_{i},e_{j})=\sum_{p=1}^{n}c_{ij}^{k}e_{k}\quad\textrm{and}\quad\mu^{\prime}(e_{i},e_{j})=\sum_{p=1}^{n}c_{ij}^{\prime k}e_{k},$
which corresponds to a point in the affine space $\mathbb{C}^{2n^{3}}$. Then a set of bilinear pairs $\mathcal{S}$ corresponds to an algebraic variety, i.e., a Zariski closed set, if there are some polynomial equations in variables $c_{ij}^{k}$, $c_{ij}^{\prime k}$ with zero locus equal to the set of structural constants of the bilinear pairs in $\mathcal{S}$. Given the identities defining a particular class of compatible $\Omega$-algebras, we can obtain a set of polynomial equations in variables $c_{ij}^{k}$, $c_{ij}^{\prime k}$. This class of $n$-dimensional compatible $\Omega$-algebras is a variety. Denote it by $\mathcal{T}_{n}$. Now, consider the following action of $\rm{GL}({\mathbb{V}})$ on ${\mathcal{T}}_{n}$
$(g*(\mu,\mu^{\prime}))(x,y):=\bigl{(}g\mu\bigl{(}g^{-1}x,g^{-1}y\bigr{)},g\mu^{\prime}\bigl{(}g^{-1}x,g^{-1}y\bigr{)}\bigr{)}$
@@
A geometric classification of a variety of algebras consists of describing the irreducible components of the variety. Recall that any affine variety can be represented as a finite union of its irreducible components uniquely. Note that describing the irreducible components of ${\mathcal{T}_{n}}$ gives us the rigid algebras of the variety, which are those bilinear pairs with an open $\textrm{GL}(\mathbb{V})$-orbit. This is due to the fact that a bilinear pair is rigid in a variety if and only if the closure of its orbit is an irreducible component of the variety. For this reason, the following notion is convenient. Denote by $\overline{{\mathcal{O}}((\mu,\mu^{\prime}))}$ the closure of the orbit of $(\mu,\mu^{\prime})\in{\mathcal{T}_{n}}$.
Definition 3.1.
@@
Furthermore, for a parametric family of algebras, we have the following notion.
Definition 3.2.
@@
Let ${{\rm T}}(*)=\{{{\rm T}}(\alpha)\colon{\alpha\in I}\}$ be a family of $n$-dimensional compatible $\Omega$-algebras of a fixed class corresponding to ${\mathcal{T}}_{n}$ and let ${{\rm T}}^{\prime}$ be another algebra. Suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure ${(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}}$. We say that the family ${{\rm T}}(*)$ degenerates to ${{\rm T}}^{\prime}$, and write ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{\{{\mathcal{O}}((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{\{{\mathcal{O}}((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$ then we call it a non-degeneration, and we write ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$.
@@
Observe that ${\rm T}^{\prime}$ corresponds to an irreducible component of $\mathcal{T}_{n}$ (more precisely, $\overline{{\rm T}^{\prime}}$ is an irreducible component) if and only if ${{\rm T}}\not\to{{\rm T}}^{\prime}$ for any $n$-dimensional compatible $\Omega$-algebra ${\rm T}$ and ${{{\rm T}(*)}\not\to{{\rm T}}^{\prime}}$ for any parametric family of $n$-dimensional compatible $\Omega$-algebras ${\rm T}(*)$. To prove a particular algebra corresponds to an irreducible component, we will use the next ideas. Firstly, since $\operatorname{dim}{\mathcal{O}}((\mu,\mu^{\prime}))=n^{2}-\operatorname{dim}\mathfrak{Der}({\rm T})$, then if ${\rm T}\to{\rm T}^{\prime}$ and ${\rm T}\not\cong{\rm T}^{\prime}$, we have that $\operatorname{dim}\mathfrak{Der}({\rm T})<\operatorname{dim}\mathfrak{Der}({\rm T}^{\prime})$, where $\mathfrak{Der}({\rm T})$ denotes the Lie algebra of derivations of ${\rm T}$. Secondly, to prove degenerations, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two compatible $\Omega$-algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\mathcal{T}}}_{n}$, respectively. Let $c_{ij}^{k}$, $c_{ij}^{\prime k}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there exist $n^{2}$ maps $a_{i}^{j}(t)\colon\mathbb{C}^{*}\to\mathbb{C}$ such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{ij}^{k}(t)$, $c_{ij}^{\prime k}(t)$ of $(\mu,\mu^{\prime})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim_{t\to 0}c_{ij}^{k}(t)=c_{ij}^{k}$ and $\lim_{t\to 0}c_{ij}^{\prime k}(t)=c_{ij}^{\prime k}$, then ${{\rm T}}\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ is called a parametrized basis for ${{\rm T}}\to{{\rm T}}^{\prime}$. In case of $E_{1}^{t},E_{2}^{t},\ldots,E_{n}^{t}$ is a parametric basis for ${\bf A}\to{\bf B}$, it will be denoted by
${\bf A}\xrightarrow{(E_{1}^{t},E_{2}^{t},\ldots,E_{n}^{t})}{\bf B}.$
Thirdly, to prove non-degenerations we may use a remark that follows from this lemma, see [3].
Lemma 3.3.
@@
Consider two compatible $\Omega$-algebras ${\rm T}$ and ${\rm T}^{\prime}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Let C be a Zariski closed in ${\mathcal{T}}_{n}$ that is stable by the action of the invertible upper ((((lower)))) triangular matrices. Then if there is a representation $(\mu,\mu^{\prime})$ of ${\rm T}$ in C, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in C.
@@
Remark 3.4.
Moreover, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two compatible $\Omega$-algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${\mathcal{T}_{n}}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Then if $\mu$, $\mu^{\prime}$, $\lambda$, $\lambda^{\prime}$ represents algebras ${\rm T}_{0}$, ${\rm T}_{1}$, ${\rm T}^{\prime}_{0}$, ${\rm T}^{\prime}_{1}$ in the affine space $\mathbb{C}^{n^{3}}$ of algebras with a single multiplication, respectively, we have ${{\rm T}_{0}\to{\rm T}^{\prime}_{0}}$ and ${\rm T}_{1}\to{\rm T}^{\prime}_{1}$. So, for example, $(0,\mu)$ can not degenerate in $(\lambda,0)$ unless $\lambda=0$.
Fourthly, to prove ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\allowbreak\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. Let $c_{ij}^{k}$, $c_{ij}^{\prime k}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there is a pair of maps $\bigl{(}f,\bigl{(}a_{i}^{j}\bigr{)}\bigr{)}$, where $f\colon\mathbb{C}^{*}\to I$ and $a_{i}^{j}\colon\mathbb{C}^{*}\to\mathbb{C}$ are such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any ${t\in\mathbb{C}^{*}}$ and the structure constants $c_{ij}^{k}(t)$, $c_{ij}^{\prime k}(t)$ of $(\mu(f(t)),\mu^{\prime}(f(t)))$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim_{t\to 0}c_{ij}^{k}(t)=c_{ij}^{k}$ and $\lim_{t\to 0}c_{ij}^{\prime k}(t)=c_{ij}^{\prime k}$, then ${{\rm T}}(*)\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ and $f(t)$ are called a parametrized basis and a parametrized index for ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, respectively. Fifthly, to prove ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$, we may use an analogous of Remark 3.4 for parametric families that follows from Lemma 3.5.
Lemma 3.5.
Consider a family of compatible $\Omega$-algebras ${\rm T}(*)$ and a compatible $\Omega$-algebra ${\rm T}^{\prime}$. Suppose ${\rm T}(*)\to{\rm T}^{\prime}$. Let C be a Zariski closed in $\mathcal{T}_{n}$ that is stable by the action of the invertible upper ((((lower)))) triangular matrices. Then if there is a representation $(\mu(\alpha),\mu^{\prime}(\alpha))$ of ${\rm T}(\alpha)$ in C for every $\alpha\in I$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in C.
@@
3.2 The geometric classification of compatible commutative
associative algebras